Circular No. 6764

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SATELLITES OF URANUS B. J. Gladman, Canadian Institute for Theoretical Astrophysics; P. D.

Nicholson and J. A. Burns, Cornell University; and J. J. Kavelaars, Mc-Master University, report the discovery of two probable distant satellites of Uranus on CCD frames obtained with the COSMIC camera on the Hale 5-m telescope at Palomar. The objects were detected by Gladman in early October on a set of twelve 360-s frames obtained on Sept. 6 and 7. The fainter object, S/1997 U 1, ∼ 6′ east of Uranus, was estimated at R = 21.9. The brighter object, S/1997 U 2, ∼ 7′ west-northwest of Uranus, was at R = 20.4—but with B ∼ 22 is fainter than the limit of the 1948 McDonald 2.1-m survey (Kuiper 1961, Planets and Satellites, p. 587); Uranus has hitherto differed from the other giant planets in that it was not known to have any distant satellites. At the request of the Central Bureau, W. Offutt, Cloudcroft, NM, began observing S/1997 U2 on Oct. 9. Both satellites were reobserved with the 5-m reflector in late October, although bad seeing permitted detection of S/1997 U 1 on only one night, when it was found ∼ 30′′ from the prediction by Nicholson on the basis of a circular uranicentric orbit. D. J. Tholen (assisted by C. Herrick and R. J. Whiteley) observed both satellites on Oct. 29 with the University of Hawaii’s 2.2-m reflector. Selected observations follow (for S/1997 U 1 and S/1997 U 2, respectively).

1997 UT α_{2000} δ_{2000} Observer
Sept. 6.21723 20^{h}31^{m}52.^{s}90 −19^{o}33′24′′.4 Gladman et al.
7.20042 20 31 45.88 −19 33 52.2  _{”}
Oct. 27.14067 20 29 34.32 −19 42 44.7  _{”}
29.29708 20 29 39.85 −19 42 26.2 Tholen
Sept. 6.22938 20^{h}30^{m}59.^{s}93 −19^{o}30′51′′.9 Nicholson et al.
7.20902 20 30 53.31 −19 31 13.3  _{”}
Oct. 9.15797 20 28 47.45 −19 37 32.8 Offutt
18.10526 20 28 48.26 −19 37 14.7  _{”}
19.15664 20 28 49.43 −19 37 08.4  _{”}
20.11488 20 28 50.71 −19 37 02.3  _{”}
22.15709 20 28 54.07 −19 36 47.1  _{”}
26.14410 20 29 03.24 −19 36 08.2 Gladman et al.
27.14091 20 29 06.04 −19 35 56.7 Offutt
27.16834 20 29 06.15 −19 35 56.5 Nicholson et al.
28.14304 20 29 09.11 −19 35 44.6  _{”}
29.28785 20 29 12.82 −19 35 29.4 Tholen
31.10198 20 29 19.30 −19 35 04.0 Offutt
1997 October 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden